The hypotenuse of an isosceles right triangle is 16 centimeters. Find the length of the two other sides of the triangle to the nearest tenth.
Answer by mananth(16946)

let each leg be x cm
.. x^2+x^2= 16^2
2x^2=256
/2
x^2= 128
x= sqrt 128
x=8sqrt2